OpenStudy (anonymous):

help! intergration by parts: when do you stop intergrating? If it is until your next intergration give you zero, then what about the special cases (e.g.: e^x) -- Thanks So much!

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Also, how do i bump?

OpenStudy (anonymous):

use until you solve it :) any example?

OpenStudy (anonymous):

e.g.: when do i stop intergrating when i do intergration by parts

OpenStudy (anonymous):

e.g. e^x is an infinite "loop"

OpenStudy (anonymous):

\[\int\limits_{ }^{ }e ^{x}dx=e ^{x} +c\]

OpenStudy (anonymous):

something like: \[\int\limits_{}^{} X^3*e^x\]

OpenStudy (anonymous):

i get (x^3)(e^x) - (3x^2)(e^x)-6x*e^x-6x*e^x-6\[\int\limits_{}^{}e^x\]

OpenStudy (anonymous):

and the last step is: -\[\int\limits_{}^{}6*e^x\]

OpenStudy (anonymous):

so isn't that an infinte loop cause i gotta intergrate until i get zero

OpenStudy (anonymous):

this will require few of those! 1) u=x^3, dv=e^x \[=x ^{3}e ^{x}-3\int\limits_{ }^{}x ^{2}e ^{x}dx=...\]

OpenStudy (anonymous):

i only know f(x) and g(x) method =(

OpenStudy (anonymous):

\[=x ^{3}e ^{x}-3x ^{2}e ^{x}+6\int\limits_{ }^{}xe ^{x}=... and one more time.. will be done!\]

OpenStudy (anonymous):

it's just a matter of names... g,f, u or v... what ever you call it.

OpenStudy (anonymous):

but when i do intergration by parts, i never stop b/c i dont know when to. If my next intergration gets to zero, i know to stop, but there are infinte loops like e^x

OpenStudy (anonymous):

i don't understand what you mean... :( can you finish your example? it's not = 0

OpenStudy (anonymous):

\[=x ^{3}e ^{x}-3x ^{2}e ^{x}+6xe ^{x}-6e ^{x}+const\]

OpenStudy (anonymous):

kk so: when you intergrate something and you have to intergrate again and again, the only way i know to stop is when i hit zero. But i never stop when i intergrate something that is a "loop" like e^x (b/c u can derive e^x for ever and still get e^x). I'm confused when to stop intergrating something.

OpenStudy (anonymous):

did you get how to do integration in your example? Do you have another one that I can see what you mean?

OpenStudy (anonymous):

yes i know the basics, but repitive intergration is where i run into trouble b/c i dont stop intergrating.

OpenStudy (anonymous):

ill look for antother example

OpenStudy (anonymous):

Normally I integrate until I find the solution. Even "loop" could be beneficial (like in sin^n & cos^n integration)...

OpenStudy (anonymous):

but how do you know when to stop? that's my problem.

OpenStudy (anonymous):

you don't... :) unless it can not be solved... some integrals could be founded only by table or computer programming... I don't think in regular Calculus program they will put something like that

OpenStudy (anonymous):

so there is no hard fast rule? just stop when it can't be solved? if so im sooooooooooo stupid

OpenStudy (anonymous):

no "fast "rules...that I know of :( I just go by the "rule", that the problems in a book & on a test - have to have solution. In real life - bummer!

OpenStudy (anonymous):

well thanks for all your help :) you rock

OpenStudy (anonymous):

you too! if you got to Calculus... you rock!

OpenStudy (anonymous):

nah u do, u can explain it :)

OpenStudy (anonymous):

thnx & good luck!!

OpenStudy (anonymous):

ty u 2